OpenStudy (anonymous):

2x + 7y / 30 i need help to evaluate this

OpenStudy (anonymous):

\[\frac{2x+7y}{30}\]?

OpenStudy (anonymous):

or \[2x+\frac{7y}{30}\]

OpenStudy (anonymous):

no the first one

OpenStudy (anonymous):

i can't solve this

OpenStudy (anonymous):

solve for x or y

OpenStudy (anonymous):

the fist sample you wrote thats my question

OpenStudy (anonymous):

will you be able to help me

OpenStudy (anonymous):

is this the full equation or is it = something???

OpenStudy (anonymous):

no just like that 2x + 7y 30

OpenStudy (anonymous):

I don't get the question

OpenStudy (anonymous):

its asking me to evaluate the equation 2x+7y 30

OpenStudy (anonymous):

I guess there is some printing error in this question,,,,,, because "Equation" means something = something and there is no = sign in your question

OpenStudy (anonymous):

the only thing that says is to evaluate

OpenStudy (anonymous):

sorry yes u right wait

OpenStudy (anonymous):

x=15 and y +3

OpenStudy (anonymous):

x=15 and y=3

OpenStudy (anonymous):

and the answer has to be as a fraction

OpenStudy (anonymous):

\[\frac{2(15)+7(3)}{30}=\frac{30+21}{30}=\frac{51}{30}=\frac{17}{10}\]

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

urwc